Please help! This is my last one and I never really done a decay problem like this...any worked out explanation would help
1/15/2016 03:00 PM 80/100 Gradebook Print Calulator Periodic Table Question 5 of 5 Incorrect A radioactive isotope decays according to the equation where N, is the amount of sample at time t, No is the origin al amount of sample, and A is the decay constant. c measurements of the remaining An unknown amount of an unidentified isotope is left to decay, and periodic measurements of the isotope are recorded in the table below Time (days) Amount of isotope (moles) f lsolope (moles) 13.0 37.0 52.0 94.0 119.0 11.9 6.54 4.5 1.6 0.84 What is A, the decay constant? Number day What is No. the original amount of radioactive material? Number mol O Previous ⓧ Give Up & View Solution Q Check Answer Next Exit Hint about us careers partners privacy policy |terms of use contact us help
Answer #1

Answer:

Decay constant : -0.050817

No = 355.3

We know that decay constant,    λ=ln(Nt/No)/t

Nt= Amount of the sample at time t

                                                  = ln(11.9/No)/13 = ln(6.54/No)/37

                                                    =    (11.9/No)/13 = (6.54/No)/37

                                                    =    (11.9/No) x 37 = (6.54/No) x 13

                                                    =    ((11.9/No) x37) - ((6.54/No)x13)

                                                    No = 440.3 - 85

                                                     No = 355.3

   To find decay constant ,

    λ= ln(0.84/355.3)/119

       = -0.050817